OpenStudy (anonymous):

how long does it take $1125 to triple if it is invested at 7% interest. compounded quarterly?

OpenStudy (anonymous):

\[3*1125=1125(1+\frac{0.07}{4})^{4t}\]

OpenStudy (anonymous):

\[0.0175^{4t}=3\]\[(4t)\log(1.0175)=\log(3)\]4t=40 t=10 years

OpenStudy (anonymous):

sorry i accumulator wrong \[t=(\frac{\log3}{\log1.0175})\div4\] t=15.83139073 years than you round